A car and a motorcycle set off from the same point to travel the same journey. The car has a start of four minutes before the motorcycle sets off. If the car travels at 80km/h and the motorcycle travels at 90km/h, how many kilometres will be travelled when the two vehicles are level?

Answer by mananth(16946)

Let the car and motorcycle meet after distance x
car speed = 80 km/h
motorcycle speed = 90 km/h
They both travel the same distance
The car takes 4 minutes more .= 4/60 = 1/15 hours
time = distance /speed
car time - motorcycle time = 1/15 hours
(x/80)-(x/90) = 1/15
LCD = 720
multiply equation by 720
9x-8x= 720/15
x= 48
They will meet after 48 km